A committee of 8 is to be formed from 16 men and 10 women.





In how many ways can the committee be formed if





i) there are no restrictions.





ii) there must be 4 men and 4 women.





iii) there should be an even number of women.





iv) there are more women than men.





v) there are atleast 6 men.

1
Expert's answer
2021-12-30T01:59:54-0500

Solution:

Given, 16 men and 10 women.

i) No. of ways when there are no restrictions

"=\ ^{26}C_8
\\=1562275"


ii) No. of ways when there must be 4 men and 4 women

"=\ ^{16}C_4\times \ ^{10}C_4
\\=382200"

iii) No. of ways when there should be an even number of women

"=\ ^{16}C_6\times \ ^{10}C_2+\ ^{16}C_4\times \ ^{10}C_4+\ ^{16}C_2\times \ ^{10}C_6+\ ^{16}C_0\times \ ^{10}C_8
\\=8008\times45+1820\times210+120\times210+1\times45
\\=767805"

iv) No. of ways when there are more women than men

"=\ ^{16}C_3\times \ ^{10}C_5+\ ^{16}C_2\times \ ^{10}C_6+\ ^{16}C_1\times \ ^{10}C_7+\ ^{16}C_0\times \ ^{10}C_8
\\=560\times 252+120\times210+16\times120+1\times45
\\=168285"

v) No. of ways when there are at least 6 men

"=\ ^{16}C_6\times \ ^{10}C_2+\ ^{16}C_7\times \ ^{10}C_1+\ ^{16}C_8\times \ ^{10}C_0
\\=8008\times45+11440\times10+12870\times1
\\=487630"